Let f(x)= x^2 – 4x –5 / 3x^2 + 2

1. Find the domain.

2.Determine the vertical asymptote(s).

3.Determine the horizontal asymptote or oblique asymptote.

4.Find the y-intercept.

5.Find the x-intercept(s)

Answer by stanbon(75887)

Let f(x)= x^2 – 4x –5 / 3x^2 + 2
1. Find the domain.
Since 3x^2+2 is never zero, the Domain is "all Real Numbers".
----------------------
2.Determine the vertical asymptote(s).
Since 3x^2+2 is never zero, there are no verticla asymptotes.
----------------------
3.Determine the horizontal asymptote or oblique asymptote.
HA: y = x^2/3x^2 = 1/3
----------------------
4.Find the y-intercept.
Let x = 0, then y = -5/2
5.Find the x-intercept(s)
Let y = 0,
Solve x^2 – 4x –5 =0
(x-5)(x+1) = 0
x = 5 or x = -1
====================
Cheers,
Stan H.